# Optimizing Calibration by Gaining Aware of Prediction Correctness

### v2

@@ Abstract
Model calibration aims to align confidence with prediction correctness. The Cross-Entropy (CE) loss is widely used for calibrator training, which enforces the model to increase confidence on the ground truth class. However, we find the CE loss has intrinsic limitations. For example, for a narrow misclassification, a calibrator trained by the CE loss often produces high confidence on the wrongly predicted class (e.g., a test sample is wrongly classified and its softmax score on the ground truth class is around 0.4), which is undesirable. In this paper, we propose a new post-hoc calibration objective derived from the aim of calibration. Intuitively, the proposed objective function asks that the calibrator decrease model confidence on wrongly predicted samples and increase confidence on correctly predicted samples. Because a sample itself has insufficient ability to indicate correctness, we use its transformed versions (e.g., rotated, greyscaled and color-jittered) during calibrator training. Trained on an in-distribution validation set and tested with isolated, individual test samples, our method achieves competitive calibration performance on both in-distribution and out-of-distribution test sets compared with the state of the art. Further, our analysis points out the difference between our method and commonly used objectives such as CE loss and mean square error loss, where the latters sometimes deviates from the calibration aim.
@@ 1 Introduction
Existing methods typically use Maximum Likelihood Estimation (MLE) to train a calibrator for the classification task, such as the Mean Square Error (MSE) loss (Tomani et al., 2022, 2023; Zhang et al., 2020) and the Cross-Entropy (CE) loss (Guo et al., 2017; Zou et al., 2023). Although these approaches demonstrate efficacy in reducing calibration errors such as expected calibration error and Brier scores, they lack theoretical guarantee that the calibration error is minimized when MLE converges. In Fig. 1, for an image which is incorrectly classified and has a relatively high probability on the ground truth class, calibrators trained by the CE or MSE loss would give high confidence on the wrongly predicted class. It means a user may trust this prediction to be true. As to be revealed in Sec. 3.4, the inherent problem of CE and MSE loss functions limits them in calibrating such test cases, but our method effectively reduces the confidence.
@@
In this paper, we derive a concrete interpretation of the goal of calibration which is then directly translated into a novel loss function that aligns with the newly interpreted goal. Specifically, we start from the general definition of calibration and its error (Guo et al., 2017) and then, under a finite test set, represent the error in a discretized form. Minimizing this discretized error gives us a very interesting and intuitive calibration goal: A correct prediction should have possibly high confidence and a wrong prediction should have possibly low confidence. Theoretically, this optimization goal can naturally reduce the overlap of confidence values between correct and incorrect predictions.
@@
We translate this goal into a loss function that enforces high confidence (i.e., 1) for correctly classified samples, and low confidence (i.e., $\frac{1}{C}$, where $C$ is the number of classes) for wrongly classified ones. We name it as the Correctness-Aware (CA) loss. Yet for this method, it is non-trivial to identify classification correctness. To address this, we propose to use transformed versions of original images as the calibrator input: consistency between their prediction results suggests prediction correctness of the original sample.
@@
Our method allows a calibrator to be trained on the in-distribution validation set and directly applied to individual test samples during inference, which is consistent with practical usage. We demonstrate the effectiveness of the proposed strategy on various test sets. In both in-distribution (IND) and out-of-distribution (OOD) test sets, our method is clearly superior to uncalibrated models and competitive compared with state-of-the-art calibrators. Moreover, our method demonstrates the potential to better separate correct and incorrect test samples using their calibrated confidence. Below we summarize the main contributions.
@@
  • Theoretically, we derive the concrete goal of model calibration that has clear semantic meaning. This allows us to design a new calibration loss function: correctly classified samples should have high confidence, while incorrectly classified ones low confidence.
@@
  • To indicate prediction correctness, we use the softmax prediction scores of transformed versions of an original image as calibrator inputs.
@@ 2 Related Work
Loss functions used in post-hoc calibration. Existing post-hoc calibrators typically use the Maximum Likelihood Estimation (MLE) for optimization (Jung et al., 2023; Tao et al., 2023). For example, Mukhoti et al. (2020) use the CE loss, while Kumar et al. (2019) use the MSE loss. Additionally, Mukhoti et al. (2020) uses the focal loss, a variant to CE, which enhances learning on wrong predictions. Guo et al. (2017) directly optimizes ECE. This paper identifies inherent problems with MLE for calibration and derives a new loss function that better aligns with calibration goal.
OOD calibration deals with distribution shifts in test sets (Tomani et al., 2021). A useful practice is modifying the calibration set to cover OOD scenarios (Tomani et al., 2021; Krishnan & Tickoo, 2020), but these methods are usually designed for specific OOD scenarios which may lead to compromised in-distribution (IND) calibration performance. Another line of works adapt to test data distribution that can be either OOD or IND, using domain adaptation (Wang et al., 2020), calibration sets from multiple domains (Gong et al., 2021), or improving calibration sets by estimating test set difficulty (Zou et al., 2023). In comparison, our calibrator is trained on the IND validation set only, and does not need test batches for update, but still demonstrates improved and competitive OOD calibration performance.
Predicting classification correctness has not been widely studied in the community. Among the few, Xia & Bouganis (2023) investigate a three-way classification problem: classify a test sample into correct prediction, wrong prediction, or an out-of-distribution sample (its category is outside the training label space). We find the task of predicting classification correctness closely connected to model calibration.
@@ 3 Approach
Figure 2: Calibration pipeline. For a given image sample $\mathbf{X}$, we first obtain its logit vector $\mathbf{z}$ and softmax vector $\mathbf{v}$. We then apply $M$ different transformations (e.g., rotation, greyscale, colorjitter, etc.) on $\mathbf{X}$ to get its transformed versions as well as their related softmax vectors as $\mathbf{v}_{i}$ ($i\in\mathcal{I}_{M}$). Indices $\mathbf{q}\in\mathbb{R}^{k}$ of the top-$k$ largest probabilities (softmax scores) in $\mathbf{v}$ are used to acquire top-$k$ scores from $\mathbf{v}_{i}$ to form the concatenated input $\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}[\mathbf{q}]$ to the calibrator. The calibrator outputs a temperature $\tau$, then being used to update the logit vector $\mathbf{z}$ to produce the calibrated softmax vector. We use our proposed Correctness-Aware (CA) loss.
@@ 3.1 Defining calibration error from its goal
Notations. We study calibration under the multi-way classification problem. Regular fonts are scalars, e.g., $\tau$; vectors are denoted by lowercase boldface letters, e.g., $\mathbf{x}$; matrices by the uppercase boldface, e.g., X for an image  [^1] For simplicity, we omit three color channels.. $\mathcal{I}_{C}$ denotes an index set of integers $\{1,...,C\}$, operator ‘;’ and $\oplus$ concatenate vectors, e.g., $\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}\!=[\mathbf{v}_{1};...;\mathbf{v}_{M}]$. A classifier $f$ takes a $d$-dimensional input $\mathbf{x}\in\mathbb{R}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}$ and its corresponding label $y\in\mathcal{I}_{C}$ with $C$ classes which are sampled from the joint distribution $p(\mathbf{x},y)=p(y|\mathbf{x})p(\mathbf{x})$. We use $\equiv$ to denote the equivalence. The output of $f$ is denoted as $f(\mathbf{X})=(\hat{y},\hat{c})$, where $\hat{y}$ and $\hat{c}$ denote the predicted class label and maximum confidence score, respectively.
Calibration goal. According to Guo et al. (2017), the goal of model calibration is to ‘align confidence with the accuracy of samples’. Based on this, existing literature define perfect calibration as:
@@
$l_{f}(\hat{c})={\|\hat{c}-\mathbb{E}^{\text{acc}}_{\hat{c}}\|}=\|\hat{c}-\int\mathbb{I}\{y_{\mathbf{x}}=\hat{y}_{\mathbf{x}}\}dp(\mathbf{x}|\hat{c})\|,$ (2)
@@ 3.2 Correctness-aware loss
However, Eq. (5) is still hard to compute because the probability density function $dp(\mathbf{x}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}|\hat{c}_{i})$ is not accessible in practice. As such, we further discretize Eq. (5), where we assume a finite number of samples $\{\mathbf{x}_{ij}\}^{m}_{j=1}$ in each distribution $p(\mathbf{x}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}|\hat{c}_{i})$. Consequently, $\mathbb{E}_{f}^{\text{emp}}$ is reformulated as:
$\begin{split}\mathbb{E}_{f}^{\text{emp}}=\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\|\hat{c_{i}}-\mathbb{I}\{y_{\mathbf{x}_{ij}}=\hat{y}_{\mathbf{x}_{ij}}\}\|.\end{split}$ (6)
Note that in practice, $m=1$, because there is only one test sample for each $p(\mathbf{x}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}|\hat{c}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i})$, i.e., a dataset only has one test sample with confidence 0.52893333We have a mild assumption that a test set has few, if any, duplicated images, where Eq. (6) will also approximately hold.. Therefore, the empirical calibration loss of classifier $f$ on a test set with $n$ samples should be:
$\mathbb{E}_{f}^{\text{emp}}=\frac{1}{n}\sum_{i=1}^{n}\|\hat{c}_{i}-\mathbb{I}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|.$ (7)
Lower and upper bounds of CA loss. Let us derive the CA loss range of Eq. (7) for a given test sample $\mathbf{x}_{i}$: (i) If the sample is correctly classified, the indicator function $\mathbb{I}^{+}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}=1$. The maximum confidence for a correctly classified sample is $\hat{c}_{i}^{+}\in(\frac{1}{C},1]$, and the loss range would be $\|\hat{c}_{i}^{+}-\mathbb{I}^{+}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|\in[0,1-\frac{1}{C})$. (ii) If the sample is wrongly classified, $\mathbb{I}^{-}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}=0$ and $\hat{c}_{i}^{-}\in(\frac{1}{C},1]$, the loss range would be $(\frac{1}{C},1]$. Let $\rho\in[0,1]$ be the ratio of correctly classified samples in $\mathcal{D}$. For the correctly classified samples, we have:
@@
$(1-\rho)\times n\times\frac{1}{C}<\sum_{i=1}^{(1-\rho)\times n}\|\hat{c}_{i}^{-}-\mathbb{I}^{-}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq(1-\rho)\times n.$ (9)
By combining Eq. (8) and (9) with Equation (7), we can now derive the lower and upper bounds of our CA loss:
$\quad\quad\!\mathbb{E}_{f}^{\text{emp}}=\frac{1}{n}\sum_{i=1}^{n}\|\hat{c}_{i}-\mathbb{I}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|$
$n\times\mathbb{E}_{f}^{\text{emp}}=\sum_{i=1}^{n}\|\hat{c}_{i}-\mathbb{I}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|$
$\quad\quad\quad\quad\!\!=\Bigg{(}\sum_{j=1}^{\rho\times n}\|\hat{c}_{j}^{+}-\mathbb{I}^{+}\{y_{\mathbf{x}_{j}}=\hat{y}_{\mathbf{x}_{j}}\}\|$
$\quad\quad\quad\quad\!\!+\sum_{k=1}^{(1-\rho)\times n}\|\hat{c}_{k}^{-}-\mathbb{I}^{-}\{y_{\mathbf{x}_{k}}=\hat{y}_{\mathbf{x}_{k}}\}\|\Bigg{)}$ (10)
$n\times\left(\frac{1-\rho}{C}\right)\leq n\times\mathbb{E}_{f}^{\text{emp}}\leq n\times\rho\times\left(1\!-\!\frac{1}{C}\right)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+n\times(1-\rho)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad=n\times\left(\frac{1-\rho}{C}+\frac{C-1}{C}\right)$
$\quad\quad\quad\quad\quad\frac{1-\rho}{C}\leq\mathbb{E}_{f}^{\text{emp}}\leq\frac{1-\rho}{C}+\frac{C-1}{C},$ (11)
@@
Theoretical insights of CA loss. First, we rewrite Eq. (7) as a sum of two CA loss components, $\mathbb{I}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}$ equals $1$ and 00 for correctly and incorrectly classified samples, respectively, and we use $\ell_{1}$ as discrepancy measurement for simplicity:
$\mathbb{E}_{f}^{\text{emp}}=\frac{1}{n}\Bigg{(}\sum_{j=1}^{\rho\times n}(1-\hat{c}_{j}^{+})\!+\!\sum_{k=1}^{(1-\rho)\times n}\hat{c}_{k}^{-}\Bigg{)},$ (12)
@@
$\mathbb{E}_{f}^{\text{emp}}=\frac{1}{n}\Bigg{(}\sum_{j=1}^{(1-\rho)\times n}(1-\hat{c}_{j}^{+})+\sum_{j=(1-\rho)n+1}^{\rho\times n}(1-\hat{c}_{j}^{+})$
$\qquad\qquad\qquad+\sum_{k=1}^{(1-\rho)\times n}\hat{c}_{k}^{-}\Bigg{)}$
$\qquad=1-\rho+\frac{1}{n}\Bigg{(}\sum_{k=1}^{(1-\rho)\times n}\hat{c}_{k}^{-}-\sum_{j=1}^{(1-\rho)\times n}\hat{c}_{j}^{+}\Bigg{)}$
$\qquad\qquad\qquad+\frac{1}{n}\sum_{j=(1-\rho)n+1}^{\rho\times n}(1-\hat{c}_{j}^{+}).$ (13)
@@
$\mathbb{E}_{f}^{\text{emp}}=(1-\rho)\mathbb{E}^{\text{diff}}+(2\rho-1)\mathbb{E}^{\text{+}}+(1-\rho)$
$\qquad\equiv(1-\rho)\mathbb{E}^{\text{diff}}+(2\rho-1)\mathbb{E}^{\text{+}}$ (14)
We notice that minimizing our CA loss is equivalent to minimizing either $\mathbb{E}^{\text{diff}}$ or $\mathbb{E}^{\text{+}}$ (omitting the constant $(1-\rho)$): (i) minimizing $\mathbb{E}^{\text{diff}}$ aims to maximize the expectation of the difference in maximum confidence scores between correct and incorrect predictions, and (ii) minimizing $\mathbb{E}^{\text{+}}$ aims to push the maximum confidence score of correctly classified samples to 1. These two objectives align well with the model calibration goal. Below we take a close look at $\mathbb{E}^{\text{diff}}$:
$\qquad\mathbb{E}^{\text{diff}}=\frac{1}{(1-\rho)\times n}\Bigg{(}\sum_{k=1}^{(1-\rho)\times n}\hat{c}_{k}^{-}-\sum_{j=1}^{(1-\rho)\times n}\hat{c}_{j}^{+}\Bigg{)}$
$=\frac{1}{1-\rho}\Bigg{(}\mathbb{E}_{f}^{\text{emp}}-\frac{1}{n}\sum_{j=(1-\rho)n+1}^{\rho\times n}\!\!\!(1-\hat{c}_{j}^{+})+\rho-1\Bigg{)}$
$\qquad\mathbb{E}_{f}^{\text{emp}}\rightarrow\frac{1-\rho}{C}\equiv\mathbb{E}^{\text{diff}}\rightarrow\frac{\Big{(}\frac{1}{C}-1\Big{)}\rho}{1-\rho}<0.$ (15)
Eq. (15) demonstrates that minimizing our CA loss $\mathbb{E}_{f}^{\text{emp}}$ during training toward the lower bound $\frac{1-\rho}{C}$ is equivalent to pushing $\mathbb{E}^{\text{diff}}$ toward $\frac{\big{(}\frac{1}{C}-1\big{)}\rho}{1-\rho}<0$. This means pushing the average maximum softmax scores of wrongly classified samples away from those of correctly classified samples, thereby reducing the overlap of confidence values between correct and incorrect predictions.

### v3

@@ Abstract
Model calibration aims to align confidence with prediction correctness. The Cross-Entropy (CE) loss is widely used for calibrator training, which enforces the model to increase confidence on the ground truth class. However, we find the CE loss has intrinsic limitations. For example, for a narrow misclassification (e.g., a test sample is wrongly classified and its softmax score on the ground truth class is 0.4), a calibrator trained by the CE loss often produces high confidence on the wrongly predicted class, which is undesirable. In this paper, we propose a new post-hoc calibration objective derived from the aim of calibration. Intuitively, the proposed objective function asks that the calibrator decrease model confidence on wrongly predicted samples and increase confidence on correctly predicted samples. Because a sample itself has insufficient ability to indicate correctness, we use its transformed versions (e.g., rotated, greyscaled, and color-jittered) during calibrator training. Trained on an in-distribution validation set and tested with isolated, individual test samples, our method achieves competitive calibration performance on both in-distribution and out-of-distribution test sets compared with the state of the art. Further, our analysis points out the difference between our method and commonly used objectives such as CE loss and Mean Square Error (MSE) loss, where the latters sometimes deviates from the calibration aim.
@@ 1 Introduction
Existing methods typically use Maximum Likelihood Estimation (MLE) to train a calibrator for the classification task, such as the Mean Square Error (MSE) loss (Tomani et al., 2022, 2023; Zhang et al., 2020) and the Cross-Entropy (CE) loss (Guo et al., 2017; Zou et al., 2023). Although these approaches demonstrate efficacy in reducing calibration errors such as Expected Calibration Error (ECE) and Brier scores, they lack theoretical guarantee that the calibration error is minimized when MLE converges. In Fig. 1, for an image which is incorrectly classified and has a relatively high probability on the ground truth class, calibrators trained by the CE or MSE loss would give high confidence on the wrongly predicted class. It means a user may trust this prediction to be true. As to be revealed in Sec. B, the inherent problem of CE and MSE loss functions limits them in calibrating such test cases.
@@
In this paper, we derive a concrete interpretation of the goal of calibration which is then directly translated into a novel loss function that aligns with the newly interpreted goal. Specifically, we start from the general definition of calibration and its error (Guo et al., 2017), and then, under a finite test set, represent the error in a discretized form. Minimizing this discretized error gives us a very interesting and intuitive calibration goal: A correct prediction should have possibly high confidence and a wrong prediction should have possibly low confidence. Theoretically, this optimization goal can naturally reduce the overlap of confidence values between correct and incorrect predictions.
@@
We translate this goal into a loss function that enforces high confidence (i.e., 1) for correctly classified samples, and low confidence (i.e., $\frac{1}{C}$, where $C$ is the number of classes) for wrongly classified ones, named Correctness-Aware (CA) loss. Nevertheless, it is non-trivial to identify classification correctness. We propose to use transformed versions of original images as the calibrator input: consistency between their prediction results suggests prediction correctness of the original sample.
@@
Our method allows a calibrator to be trained on the in-distribution validation set and directly applied to individual test samples during inference. We demonstrate the effectiveness of the proposed strategy on various test sets. In both in-distribution (IND) and out-of-distribution (OOD) test sets, our method is clearly superior to uncalibrated models and competitive compared with state-of-the-art calibrators. Moreover, our method demonstrates the potential to better separate correct and incorrect test samples using their calibrated confidence. Below we summarize the main contributions.
@@
  • Theoretically, we derive the concrete goal of model calibration that has a clear semantic meaning. This allows us to design a new calibration loss function: correctly classified samples should have high confidence, while incorrectly classified ones with low confidence.
@@
  • To indicate prediction correctness, we use the softmax prediction scores of transformed versions of the original image as calibrator inputs.
@@ 2 Related Work
Loss functions used in post-hoc calibration. Existing post-hoc calibrators typically use the Maximum Likelihood Estimation (MLE) for optimization (Jung et al., 2023; Tao et al., 2023). For example, Mukhoti et al. (2020) use the CE loss, while Kumar et al. (2019) use the MSE loss. Additionally, Mukhoti et al. (2020) uses the focal loss, a variant to CE, which enhances learning on wrong predictions. Guo et al. (2017) directly optimizes the Expected Calibration Error (ECE). This paper identifies inherent problems with MLE for calibration and derives a new loss function that better aligns with calibration goal.
OOD calibration deals with distribution shifts in test sets (Tomani et al., 2021). A useful practice is modifying the calibration set to cover OOD scenarios (Tomani et al., 2021; Krishnan & Tickoo, 2020), but these methods are usually designed for specific OOD scenarios which may lead to compromised IND calibration performance. Another approach adapts to test data distribution, whether OOD or IND, using domain adaptation (Wang et al., 2020), calibration sets from multiple domains (Gong et al., 2021), or improving calibration sets by estimating test set difficulty (Zou et al., 2023). In comparison, our calibrator is trained on the IND validation set only, and does not need test batches for update, but still demonstrates improved and competitive OOD calibration performance.
Predicting classification correctness has not been widely studied. Among the few, Xia & Bouganis (2023) investigate a three-way classification problem: classify a test sample into correct prediction, wrong prediction, or an out-of-distribution sample (its category is outside the training label space). We find the task of predicting classification correctness closely connected to model calibration.
@@ 3.1 Defining calibration error from its goal
Notations. We study calibration under the multi-way classification problem. Regular fonts are scalars, e.g., $\tau$; vectors are denoted by lowercase boldface letters, e.g., $\mathbf{x}$; matrices by the uppercase boldface, e.g., X for an image. [^1] For simplicity, we omit three color channels. $\mathcal{I}_{C}$ denotes an index set of integers $\{1,...,C\}$, operator ‘;’ and $\oplus$ concatenate vectors, e.g., $\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}\!=\![\mathbf{v}_{1};...;\mathbf{v}_{M}]$. A classifier $f$ takes a $d$-dimensional input $\mathbf{x}\in\mathbb{R}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}$ and its corresponding label $y\!\in\!\mathcal{I}_{C}$ with $C$ classes which are sampled from the joint distribution $p(\mathbf{x},y)\!=\!p(y|\mathbf{x})p(\mathbf{x})$. We use $\equiv$ to denote the equivalence. The output of $f$ is denoted as $f(\mathbf{X})\!=\!(\hat{y},\hat{c})$, where $\hat{y}$ and $\hat{c}$ denote the predicted class and maximum confidence score, respectively.
Calibration goal. According to Guo et al. (2017), the goal of model calibration is to “align confidence with the accuracy of samples.” Based on this, existing literature define perfect calibration as:
@@
$l_{f}(\hat{c})=D(\hat{c}\|\mathbb{E}^{\text{acc}}_{\hat{c}})={\|\hat{c}-\mathbb{E}^{\text{acc}}_{\hat{c}}\|}$
$\quad\quad=\|\hat{c}-\int\mathbb{I}\{y_{\mathbf{x}}=\hat{y}_{\mathbf{x}}\}dp(\mathbf{x}|\hat{c})\|,$ (2)
@@ 3.2 Correctness-aware loss
However, Eq. (5) is still hard to compute because the probability density function $dp(\mathbf{x}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}|\hat{c}_{i})$ is not accessible in practice. As such, we further discretize Eq. (5), where we assume a finite number of samples $\{\mathbf{x}_{ij}\}^{m}_{j=1}$ in each distribution $p(\mathbf{x}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}|\hat{c}_{i})$. Consequently, $\mathbb{E}_{f}^{\text{emp}}$ is approximated as:
$\begin{split}\mathbb{E}_{f}^{\text{emp}}\approx\frac{1}{n}\sum_{i=1}^{n}\|\hat{c_{i}}-\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}\{y_{\mathbf{x}_{ij}}=\hat{y}_{\mathbf{x}_{ij}}\}\|.\end{split}$ (6)
Note that in practice, $m=1$, because there is only one test sample for each $p(\mathbf{x}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}|\hat{c}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i})$, e.g., a dataset only has one test sample with confidence 0.52893.333We have a mild assumption that a test set has few, if any, duplicated images, where Eq. (6) will also approximately hold. Therefore, we define the correctness aware (CA) loss as the empirical calibration loss of classifier $f$ on a test set with $n$ samples:
$\mathcal{L}_{\text{CA}}=\frac{1}{n}\sum_{i=1}^{n}\|\hat{c}_{i}-\mathbb{I}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|.$ (7)
Figure 2: Calibration pipeline. For a given image sample $\mathbf{X}$, we first obtain its logit vector $\mathbf{z}$ and softmax vector $\mathbf{v}$. We then apply $M$ different transformations (e.g., rotation, greyscale, colorjitter, etc.) on $\mathbf{X}$ to get its transformed versions as well as their related softmax vectors as $\mathbf{v}_{i}$ ($i\in\mathcal{I}_{M}$). Indices $\mathbf{q}\in\mathbb{R}^{k}$ of the top-$k$ largest probabilities (softmax scores) in $\mathbf{v}$ are used to acquire top-$k$ scores from $\mathbf{v}_{i}$ to form the concatenated input $\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}[\mathbf{q}]$ to the calibrator. The calibrator outputs a temperature $\tau$, then being used to update the logit vector $\mathbf{z}$ to produce the calibrated softmax vector. We use our proposed Correctness-Aware (CA) loss (Sec. 3.2).
Lower and upper bounds of CA loss. Let us derive the CA loss range of Eq. (7) for a given test sample $\mathbf{x}_{i}$: (i) If the sample is correctly classified, the indicator function $\mathbb{I}^{+}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}=1$. The maximum confidence range for a correctly classified sample is $\hat{c}_{i}^{+}\in(\frac{1}{C},1]$, and the loss range would be $\|\hat{c}_{i}^{+}-\mathbb{I}^{+}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|\in[0,1-\frac{1}{C})$. (ii) If the sample is wrongly classified, $\mathbb{I}^{-}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}=0$ and $\hat{c}_{i}^{-}\in(\frac{1}{C},1]$, the loss range would be $(\frac{1}{C},1]$. Let $\rho\in[0,1]$ be the ratio of correctly classified samples in $\mathcal{D}$. For the correctly classified samples, we have:
@@
$(1-\rho)\!\times\!n\!\times\!\frac{1}{C}\!<\!\!\!\!\!\sum_{i=1}^{(1-\rho)\times n}\!\!\!\!\!\|\hat{c}_{i}^{-}\!-\!\mathbb{I}^{-}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}\|\!\leq\!(1\!-\!\rho)\!\times\!n.$ (9)
Combining Eq. (8) and (9) with Equation (7), we can derive the lower and upper bounds of CA loss:
$n\!\times\!\mathcal{L}_{\text{CA}}=\sum_{j=1}^{\rho\times n}\|\hat{c}_{j}^{+}\!-\!\mathbb{I}^{+}\{y_{\mathbf{x}_{j}}\!=\!\hat{y}_{\mathbf{x}_{j}}\}\|$
$\quad\quad\quad\quad+\!\!\sum_{k=1}^{(1-\rho)\times n}\|\hat{c}_{k}^{-}-\mathbb{I}^{-}\{y_{\mathbf{x}_{k}}\!=\!\hat{y}_{\mathbf{x}_{k}}\}\|,$ (10)
$n\!\times\!\left(\!\frac{1-\rho}{C}\!\right)\leq n\!\times\!\mathcal{L}_{\text{CA}}\leq n\times\left(\frac{1-\rho}{C}+\frac{C-1}{C}\right),$
$\quad\quad\quad\frac{1-\rho}{C}\leq\mathcal{L}_{\text{CA}}\leq\frac{1-\rho}{C}+\frac{C-1}{C},$ (11)
@@
Theoretical insights of CA loss. First, we rewrite Eq. (7) as a sum of two CA loss components, $\mathbb{I}\{y_{\mathbf{x}_{i}}=\hat{y}_{\mathbf{x}_{i}}\}$ equals $1$ and 00 for correctly and incorrectly classified samples, respectively, and we use $\ell_{1}$-distance for simplicity:
$\mathcal{L}_{\text{CA}}=\frac{1}{n}\Bigg{(}\sum_{j=1}^{\rho\times n}(1-\hat{c}_{j}^{+})\!+\!\sum_{k=1}^{(1-\rho)\times n}\hat{c}_{k}^{-}\Bigg{)},$ (12)
@@
$\mathcal{L}_{\text{CA}}\!=\!\frac{1}{n}\Bigg{(}\sum_{j=1}^{\!\!(1\!-\!\rho)\!\times\!n}(1\!-\!\hat{c}_{j}^{+})+\!\!\!\!\sum_{j=(1-\rho)n+1}^{\rho\times n}\!\!\!\!(1\!-\!\hat{c}_{j}^{+})+\!\!\!\sum_{k=1}^{(1-\rho)\times n}\!\!\hat{c}_{k}^{-}\Bigg{)}$
$=1\!-\!\rho\!+\!\frac{1}{n}\Bigg{(}\!\!\sum_{k=1}^{(1-\rho)\times n}\!\!\!\hat{c}_{k}^{-}\!-\!\!\!\!\sum_{j=1}^{(1-\rho)\!\times\!n}\!\!\!\hat{c}_{j}^{+}\!\!\Bigg{)}\!+\!\frac{1}{n}\!\!\!\!\sum_{j=(1-\rho)n+1}^{\rho\times n}\!\!\!\!(1\!-\!\hat{c}_{j}^{+}).$ (13)
@@
$\mathcal{L}_{\text{CA}}\!\!=\!\!(1\!-\!\rho)\mathbb{E}^{\text{diff}}\!\!+\!\!(2\rho\!-\!1)\mathbb{E}^{\text{+}}\!+\!(1\!-\!\rho).$ (14)
We notice that minimizing our CA loss is equivalent to minimizing either $\mathbb{E}^{\text{diff}}$ or $\mathbb{E}^{\text{+}}$ (omitting the constant $(1-\rho)$): (i) minimizing $\mathbb{E}^{\text{diff}}$ aims to maximize the expectation of the difference in maximum confidence scores between correct and incorrect predictions, and (ii) minimizing $\mathbb{E}^{\text{+}}$ aims to push the maximum confidence score of correctly classified samples to 1. These two objectives align well with the model calibration goal (Sec. 3.1). Below we take a close look at $\mathbb{E}^{\text{diff}}$:
$\mathbb{E}^{\text{diff}}\!=\!\frac{1}{(1\!-\!\rho)\!\times\!n}\Bigg{(}\!\sum_{k=1}^{(1-\rho)\times n}\hat{c}_{k}^{-}\!-\!\sum_{j=1}^{(1-\rho)\times n}\hat{c}_{j}^{+}\Bigg{)}$
$\quad\quad\!\!\!=\frac{1}{1-\rho}\Bigg{(}\mathcal{L}_{\text{CA}}-\frac{1}{n}\sum_{j=(1-\rho)n+1}^{\rho\times n}\!\!\!(1-\hat{c}_{j}^{+})+\rho-1\Bigg{)},$
$\mathcal{L}_{\text{CA}}\rightarrow\frac{1-\rho}{C}\equiv\mathbb{E}^{\text{diff}}\rightarrow\frac{\Big{(}\frac{1}{C}-1\Big{)}\rho}{1-\rho}<0.$ (15)
Eq. (15) demonstrates that minimizing our CA loss $\mathcal{L}_{\text{CA}}$ during training toward the lower bound $\frac{1-\rho}{C}$ is equivalent to pushing $\mathbb{E}^{\text{diff}}$ toward $\frac{\big{(}\frac{1}{C}-1\big{)}\rho}{1-\rho}<0$. This means pushing the average maximum softmax scores of wrongly classified samples away from those of correctly classified samples, thereby reducing the overlap of confidence values between correct and incorrect predictions.